OpenStudy (anonymous):

what is the value of x -3x^2+18x-24=0

OpenStudy (anonymous):

do you know factoring or the quadratic formula ? which method you wanna use?

OpenStudy (anonymous):

quadratic formula

OpenStudy (anonymous):

ok... let's make things easier so we can work with smaller numbers.. dividing both sides by -3 we should get: x^2 - 6x + 8 = 0 can you identify a, b, c ?

OpenStudy (anonymous):

\[\begin{array}{*{20}c} {x = \frac{{ - b \pm \sqrt {b^2 - 4ac} }}{{2a}}} & {{\rm{when}}} & {ax^2 + bx + c = 0} \\ \end{array}\] Your equation gives us: \[x^2-6x+8=0\]